Math 544 (Section 001)  University of South Carolina
D. Meade  Fall 1996

Homework Problems for Chapter 2 The Vector Space R^{n}

Section Page Problems
2.1  120 # 13, 14^{∗}, 15, 18, 21, 23, 24, 26, 29^{∗}
2.2  127 # 1, 2, 4, 5, 7^{∗}, 9, 10, 12, 13, 14, 15, 18^{∗}, 20, 29, 30
2.3  139 # 5, 6, 15, 16^{∗}, 17, 21, 22, 27, 28, 31^{∗}, 34, 40, 42
2.4  153 # 1, 2, 6, 9, 12, 13^{∗}, 19, 21, 25, 26, 27^{∗}, 32, 33
2.5  166 # 2, 4, 5, 6, 9^{∗}, 15, 16, 19, 22^{∗}, 25, 28, 32, 35
2.6  177 # 3^{∗}, 5, 8, 9, 10, 13, 14^{∗}, 15
2.7  193 # 2, 3^{∗}, 6, 7, 9, 10, 11, 19^{∗}, 22, 25, 26, 32, 40bc
2.8  204 # 1, 4^{∗}, 11, 13, 19^{∗}
2.9  213 # 1^{∗}, 2, 4, 6, 7, 8^{∗}, 13
Review 214 Supplemental Computational Exercises

# 1–9, 11, 12–15

Review 216 Supplemental Conceptual Exercises

# 1–12, 13–18

\underline{Notes}

• Problems marked with an asterisk (^{∗}) will be collected and graded. In general, problems will be due the day following the lecture in which the section was discussed.

• The selected problems require a mixture of computation and theory to answer. In most cases the problems are not difficult once the question is identified and understood. The key to this is a thorough understanding of the vocabulary and various connections between the concepts.

• If you need more practice, there are many other problems in the text that are similar to the ones I have assigned.

• There are two sections of review problems at the end of each chapter. None of these problems will be collected for grading. These problems should be a good test of your understanding of the main points in the chapter. I suggest solving these problems as one part of your preparation for an exam.

• Be sure you know how to do all of these problems. Do not hesitate to ask questions – in class, by e-mail, or during office hours.